# High-Resolution Flood Probability Mapping Using Generative Machine Learning with Large-Scale Synthetic Precipitation and Inundation Data

### v1

@@ 3.3 Synthetic Rainfall Event Assessment
Using the strategic filtering and sampling methods introduced in Section 2.5, we generated 10,000 rainfall events through parallel computing. For each event, the global duration was sampled from the tailored pool. As the number of event simulations increases, the overall synthetic precipitation events tend to closely resemble true events. We used 20-nearest neighbors (20-NN) to fill the empty cells after stringent filtering, using averaged precipitation-based feature values. It is common to observe a pepper-salt appearance in cumulative and peak precipitation maps due to the independence of point data in CTGAN records generation. To address this, we applied 20-NN smoothing for synthetic precipitation events with duration of 2 hours or less, and 50-NN smoothing for longer duration, to enhance the realism of the synthetic events. From these smoothed cumulative and peak precipitation features, we derived heavy cumulative and peak precipitation ratios, resulting in 21 input features for the selected depth estimator, MaxFloodCast V2, to generate synthetic inundation depths and complete the precipitation-flood event. The event generation process took an average of 10 minutes to generate 100 rainfall events in parallel, consuming 100 GB of memory on an AMD EPYC 7702P 64-core processor.
Figure 5: Synthetic flood event with 15 hours global precipitation. The synthetic cumulative precipitation and peak precipitation are processed by 50-NN smoother. Maps generated by Geopandas Python package.
Fig. 5 illustrates a synthetic precipitation-flood event with a global duration of 15 hours, comprising four subplots: synthetic cumulative precipitation, synthetic peak precipitation, precipitation duration, and synthetic inundation depth. The cumulative and peak precipitation maps show significant spatial variability, with higher values in central and northern regions. Despite setting a global duration of 15 hours, the method effectively creates localized precipitation patterns. The inundation depth map highlights flood-prone regions, particularly along river channels and low-lying areas. This synthetic event realistically represents potential flood scenarios, enhancing flood risk assessment and mitigation efforts.
Figure 6: Comparison of Sampled Training Depth Distribution with Downsampled Synthetic Depth Distribution.
@@
Mean 2.1171 0.8634 0.5578 0.8280
Std 1.4817 0.0554 0.0981 0.4575
Min 0.3184 0.3448 0.0010 0.0231
25% 1.0865 0.8398 0.5193 0.5430
50% 1.5940 0.8730 0.5663 0.7616
75% 2.7402 0.8993 0.6115 1.0106
Max 16.850 0.9861 0.8923 6.2082
@@
We generated 10,000 synthetic precipitation-flood events, providing each cell with a comprehensive synthetic depth distribution for direct comparison with the training dataset. To assess the performance of the synthetic events, we utilized four key metrics—RMSE, Cosine Similarity, Correlation, and KL Divergence—comparing them against the limited number of sampled training events. To ensure a fair comparison, the synthetic events were downsampled to match the size of the training data. This downsampling process was repeated 50 times to minimize variability introduced by randomness. A detailed statistical comparison between the normalized versions of the training and synthetic datasets is provided in Table 4, with the corresponding visualizations shown in Fig. 6.
@@
The synthetic data demonstrates strong overall performance, as evidenced by the RMSE and cosine similarity metrics, indicating that the generated events align closely with the training data and accurately capture key hydrological characteristics and trends. Notably, the mean RMSE for non-channel cells is 2.12 ft, substantially lower than the 5.96 ft recorded for channel cells. This disparity suggests that the synthetic data more effectively represents regions with simpler hydrodynamics, while the higher RMSE in channel cells reflects the inherent difficulty in modeling real-world depth distributions in areas characterized by dynamic flow patterns and greater variability. Nevertheless, the synthetic data performs well even in these more complex regions, achieving a mean cosine similarity of 0.85 for channel cells, highlighting the model’s ability to capture the overall structure and trends of the real data.
@@
In non-channel cells, the model excels with a mean cosine similarity of 0.86 and a relatively low KL divergence of 0.83. These metrics suggest that the synthetic data closely matches the probability distributions of the real data in these regions, where simpler hydrological behavior makes depth distributions easier to replicate. In contrast, the KL divergence for channel cells, though still low at 0.80, indicates some deviation between the synthetic and real data, which is expected given the complexity and variability of flood behavior in these more dynamic areas.
@@
The correlation between the synthetic and real datasets further reinforces the model’s effectiveness. Non-channel cells exhibit a mean correlation of 0.56, while channel cells show a slightly lower but still robust correlation of 0.55. This highlights the model’s ability to capture linear relationships across both cell types, though the more intricate hydrodynamic processes in channel cells present additional challenges. Hydrologically, these findings are particularly significant. Channel cells, which are more prone to flooding due to their dynamic flow patterns, underscore the model’s strength in capturing intricate interactions even under complex conditions. The model’s ability to generalize across non-channel cells, reflected in their lower RMSE and KL divergence values, is equally noteworthy. It demonstrates the model’s capacity to explore a variety of flood scenarios without overfitting to specific events.
@@
Building on this analysis, we generated flood probability maps. Fig. 7 presents synthetic flood probability maps for various inundation depths: 1 ft, 2 ft, 4 ft, and 6 ft. These maps highlight the probability of exceeding specified flood depths across the study area. The probability values specify the proportion of precipitation-flood events that cause flood depths of certain level or higher in a given area across the 10,000 precipitation-flood events. These values capture the flood risk of certain areas for certain inundation levels. A high $P(\text{depth}\geq 6ft)$ in a given area means that for a generated generic flood event there is a high probability that flood level will equal or exceed 6 ft. In other words, the probability value serves as a quantification of risk to a certain depth of flooding. This interpretation should be distinct from annual flood probability which is used in standard flood mapping methods. As shown in Fig. 7, at 1 ft and 2 ft depths, flood probabilities greater than 0.5 (high probability) are predominantly observed in the northeastern, southeastern, and central regions of the study area, while lower probabilities are seen in the western and southern areas. As the depth threshold increases to 4 ft and 6 ft, the high-probability zones become more concentrated along rivers/channels and low-lying regions, while most of the area shows low flood risk.
(a) $P(\text{depth}\geq 1\text{ft})$
(b) $P(\text{depth}\geq 2\text{ft})$
(c) $P(\text{depth}\geq 4\text{ft})$
(d) $P(\text{depth}\geq 6\text{ft})$
@@ 4 Discussion
In this study, we presented an innovative methodology, Flood-Precip GAN, for generating high-resolution flood probability maps using synthetic precipitation-flood events. Our approach leveraged advanced machine learning techniques, including Conditional Generative Adversarial Networks (CTGAN) [23] as point location and precipitation-based features generator and XGBoost-based models MaxFloodCast [20] as synthetic depth estimator, to overcome the limitations of traditional flood modeling that rely heavily on historical data. To address the limitations in feature dimensionality found in existing CTGAN tools and the challenge of generating individual precipitation-flood events from a massive synthetic point cloud, we designed a robust and systematic workflow. This approach includes cell-level pool creation, strategic sampling, noise smoothing, depth synthesis, and the eventual formation of individual events. Each step in the process is carefully structured to avoid overfitting and ensure that the generated synthetic events accurately reflect real-world flood scenarios while maintaining computational efficiency and high fidelity to hydrological patterns. By generating 10,000 synthetic precipitation-flood events, we created a comprehensive dataset that captures a wide spectrum of potential flood scenarios, significantly improving the robustness of flood risk assessments. Based on the metrics between the training and synthetic events across overall, channel, and non-channel aspects shown on Table 4, the overall Cosine Similarity, with a mean of 0.86 and a standard deviation of 0.05, indicates a strong alignment between the synthetic and training data. The RMSE values highlight a notable difference between channel and non-channel cells, reflecting the challenges in accurately modeling flood depth distribution in regions with dynamic flows and greater variability. These findings demonstrate the model’s ability to capture key trends and linear relationships, particularly in hydrologically complex regions such as channel cells, which are more susceptible to flooding. Additionally, the synthetic data’s capacity to generalize beyond the training events highlights its potential for producing realistic flood simulations, which is critical for accurate risk assessment and effective mitigation planning.
@@
The flood probability maps provide an in-depth spatial analysis of flood risk, revealing critical high-risk zones, particularly along river channels and low-lying areas. These maps emphasize the importance of depth-specific flood management strategies, as the distinction between channel and non-channel cells highlights the need to account for topographical and hydrological variability. This is particularly crucial in regions where intricate flow dynamics, terrain elevation, and slope significantly influence flood behavior. The high resolution of the maps also allows for practical applications, such as assessing the flood risk of buildings by integrating lowest floor elevation datasets, identifying structures at risk of inundation over 2 ft. Additionally, these maps can evaluate transportation network vulnerability, as roads in areas with a high probability of exceeding 2 ft of inundation would become impassable, disrupting transportation and emergency response efforts. The non-binary nature of the maps further allows for comparative flood risk assessments across different inundation levels, enabling more nuanced and targeted flood mitigation strategies across diverse landscapes.
Moreover, the synthetic dataset’s ability to simulate a broad range of generic flood scenarios, combined with its strong alignment to the training data, reinforces its applicability for future flood risk assessments. These findings suggest that this approach could play a pivotal role in enhancing predictive capabilities for flood events, enabling more effective planning and resource allocation for flood management in both simple and hydrologically complex regions.
Despite these advancements and novelties, the current approach has limitations in generating local precipitation events, even if the synthesizer has capability to generate records with no precipitation under a global precipitation duration. Plus, the computational complexity is significant, taking approximately 10 minutes and 100GB CPU memory to generate 100 precipitation-flood events. Future work should focus on improving the efficiency of the computational process and enhancing the model’s ability to generate more localized precipitation patterns.
@@
Overall, this study demonstrates the efficacy of integrating machine learning and generative models to produce realistic and comprehensive synthetic precipitation-flood events for producing flood probability maps. The synthetic flood probability maps generated in this research provide valuable insights for urban planners, emergency management professionals, and policymakers, enabling more effective flood preparedness and mitigation strategies. This work represents a significant step forward in flood risk management, offering a scalable and flexible solution to address the limitations of current flood risk mapping and characterization approaches. This study also contributes to efforts to expand simulated (synthetic) data for AI model development for disaster resilience use cases. Both real and synthetic (simulated) data are crucial for the development of robust AI models for various disaster resilience use cases ranging from risk prediction to situational awareness and impact assessment. Future studies can leverage GAN and other generative ML methods for augmenting limited real disaster datasets and produce datasets suitable for training and testing novel ML models for various disaster resilience applications.
Acknowledgements
@@
The authors would like to acknowledge funding support from the National Science Foundation under CRISP 2.0 Type 2, grant 1832662, and the Texas A&M X-Grant Presidential Excellence Fund. Any opinions, findings, conclusions, or recommendations expressed in this research are those of the authors and do not necessarily reflect the view of the funding agencies.

### v2

@@ 3.3 Synthetic Rainfall Event Assessment
Using the strategic filtering and sampling methods outlined in Section 2.5, we generated 10,000 synthetic rainfall events through parallel computing. For each event, the sampled global duration level guided and constrained the selection of precipitation features from each cell’s tailored pool. Leveraging the 27-element precipitation level combination pool described in Section 2.5, the cell-level sampling preserved regular precipitation trends while maintaining the probability of extreme scenarios. As the number of simulated events increased, the overall synthetic precipitation patterns closely aligned with those in the training dataset. Due to the independence and nonlinearity of point data in CTGAN-generated records, cumulative and peak precipitation maps often exhibited a "pepper-salt" appearance. To mitigate this, we applied KNN smoothing, selecting the value of $K$ based on the synthetic event’s duration following that the longer duration correspond to larger $K$ values, and vice versus. From these smoothed precipitation-based features, we derived HCPR and HPPR, resulting in 21 input features for the depth estimator, MaxFloodCast V2, which was then used to generate synthetic inundation depths and complete the precipitation-flood event. The event generation process averaged 10 minutes for 100 rainfall events in parallel, utilizing 100 GB of memory on an AMD EPYC 7702P 64-core processor.
Figure 4: Synthetic flood event with 15 hours global precipitation. All the attribute maps are in the same scale and they share the same color bar with different units: inch, inch, hour and feet respectively. The synthetic cumulative precipitation and peak precipitation are processed by 50-NN smoother. Maps generated by Geopandas Python package.
Fig. 4 illustrates a synthetic precipitation-flood event with a global duration of 15 hours, comprising four subplots: synthetic cumulative precipitation, synthetic peak precipitation, precipitation duration, and synthetic inundation depth. In Harris County this 15-hour storm event would have a return period of approximately 200 years 22. Cumulative and peak precipitation maps show some spatial variability, as expected for a relatively long but still sub-daily rainfall. The inundation depth map highlights flood-prone regions, particularly along river channels and low-lying areas. This result is consistent with the expectation that a 200-year storm would cause multiple floodings at the local level in the area.
Figure 5: Histograms of Flood Assessment Metrics: Comparison of metrics between channel cells and non-channel cells, based on the differences between the sampled training depth distribution and the downsampled synthetic depth distribution.
@@
Mean 2.0992 0.8696 0.5525 0.8224
Std 1.4678 0.0562 0.1013 0.4532
Min 0.3221 0.3511 0.0010 0.0240
25% 1.0912 0.8415 0.5240 0.5521
50% 1.6098 0.8772 0.5644 0.7559
75% 2.7623 0.9028 0.6183 1.0194
Max 16.940 0.9815 0.8897 6.2550
3.4 Synthetic Rainfall Distribution Assessment
@@
We generated 10,000 synthetic precipitation-flood events, each cell having a comprehensive synthetic depth distribution for direct comparison with the training dataset. To assess the performance of the synthetic events quantitatively, we utilized four key metrics—RMSE, Cosine Similarity, Correlation, and KL Divergence—comparing them against the limited number of sampled training events. To ensure a fair comparison, the synthetic events were downsampled to match the size of the training data. This downsampling process was repeated 50 times to minimize variability introduced by randomness. A detailed statistical comparison between the normalized versions of the training and the synthetic datasets is provided in Table 4, with the corresponding visualizations shown in Fig. 5.
@@
The sampled synthetic flood depth distribution demonstrates comparable overall performance, as evidenced by the RMSE and cosine similarity metrics, indicating that the generated events align closely with the training data and accurately capture key hydrological characteristics and trends. Notably, the mean RMSE for non-channel cells is 2.10 ft, substantially lower than the 5.99 ft recorded for channel cells. This disparity suggests that the synthetic data more effectively represents regions with simpler hydrodynamics, while the higher RMSE in channel cells reflects the inherent difficulty in modeling real-world depth distributions in areas characterized by dynamic flow patterns and greater variability. Nevertheless, the synthetic data performs well even in these more complex regions, achieving a mean cosine similarity of 0.85 for channel cells, highlighting the model’s ability to capture the overall structure and trends of the real data.
@@
In non-channel cells, the model excels with a mean cosine similarity of 0.87 and a relatively low KL divergence of 0.82. These metrics suggest that the synthetic data closely matches the probability distributions of the real data in these regions, where simpler hydrological behavior makes depth distributions easier to replicate. In contrast, the KL divergence for channel cells, though still low at 0.78, indicates some deviation between the synthetic and real data, which is expected given the complexity and variability of flood behavior in these more dynamic areas.
@@
The correlation between the synthetic and real datasets further reinforces the model’s effectiveness. Non-channel cells exhibit a mean correlation of 0.55, while channel cells show a slightly lower but still robust correlation of 0.54. This highlights the model’s ability to capture linear relationships across both cell types, though the more intricate hydrodynamic processes in channel cells present additional challenges. Hydrologically, these findings are particularly significant. Channel cells, which are more prone to flooding due to their dynamic flow patterns, underscore the model’s strength in capturing intricate interactions even under complex conditions. The model’s ability to generalize across non-channel cells, reflected in their lower RMSE and KL divergence values, is equally noteworthy. It demonstrates the model’s capacity to explore a variety of flood scenarios without overfitting to specific events.
@@
3.5 Flood probability maps
Flood depths generated by a large number of synthetic precipitation-flood events represent a foundational data set from which one can estimate flood probability. Probability is calculated cell-wise as the proportion of synthetic events that cause a certain value of flood depth or higher. Using the 10,000 synthetic events simulated in Section 3.3, we built synthetic flood probability maps for the following inundation depths: 1 ft, 2 ft, 4 ft, and 6 ft (Fig. 6). In Fig. 6.IV, for example, a high $P(\text{depth}\geq 6ft)$ in a cell means that for a generic flood event there is a high probability that flood level will equal or exceed 6 ft in that cell. In other words, the probability value serves as a quantification of risk for a certain flooding depth. This interpretation should be distinct from annual flood probability which is traditionally used in flood mapping. As shown in Fig. 6, at 1 ft and 2 ft depths, flood probabilities greater than 0.5 (high probability) are predominantly observed in the northeastern, southeastern, and central regions of the study area, while lower probabilities are seen in the western and southern areas. As the depth threshold increases to 4 ft and 6 ft, the high-probability zones become more concentrated along rivers/channels and low-lying regions.
(I) $P(\text{depth}\geq 1\text{ft})$
(II) $P(\text{depth}\geq 2\text{ft})$
(III) $P(\text{depth}\geq 4\text{ft})$
(IV) $P(\text{depth}\geq 6\text{ft})$
@@ 4 DISCUSSION
In this study, we presented an innovative application methodology, Precipitation-Flood Depth Generative Pipeline, for generating high-resolution flood probability maps using synthetic precipitation-flood events. Our approach leveraged advanced machine learning techniques, including Conditional Generative Adversarial Networks (CTGAN) 32 as point location, and precipitation-based features generator and XGBoost-based models MaxFloodCast 15 as synthetic depth estimator, to overcome the limitations of traditional flood modeling that rely on historical data and records, not always available. To address the limitations in feature dimensionality found in existing CTGAN tools and the challenge of generating individual precipitation-flood events from a massive synthetic point cloud, we designed a robust and systematic workflow. This includes cell-level pool creation, strategic sampling, noise smoothing, depth synthesis, and the eventual formation of individual events. Each step in the process is carefully structured to avoid overfitting and ensure that the generated synthetic events accurately reflect real-world flood scenarios while maintaining computational efficiency and high fidelity to hydrological patterns. By generating 10,000 synthetic precipitation-flood events, we created a comprehensive dataset that captures a wide spectrum of potential flood scenarios, increasing the robustness of flood risk assessments. The RMSE values highlight a notable difference between channel and non-channel cells, reflecting the challenges in accurately modeling flood depth distribution in regions with dynamic flows and greater variability. These findings demonstrate the model’s ability to capture key trends and non-linear relationships, particularly along channels and in floodplain areas, which are more susceptible to flooding. Additionally, the synthetic data’s capacity to generalize beyond the training events highlights its potential for producing possible flood scenarios, which is critical for accurate risk assessment and mitigation planning.
4.1 Spatial nature of weather and flooding
Despite the advancements introduced with the Precipitation-Flood Depth Generative Pipeline, this approach has some limitations in generating local precipitation events. While CTGAN can generate records with no precipitation, we observed a tendency to generate storms that affect the majority of the study area. The prevalence of large-extent storms in the synthetic dataset influenced flood depth distributions at the cell level and consequently the flood probability maps. Flood probability maps currently in use in the study area show the extent of the 100-year and 500-year floodplain, but do not provide information on associated flood depths at the largest scale, making difficult a direct comparison with the synthetic flood probability maps illustrated in Fig. 6. However, the flood probability for low depth values (1 and 2 feet) is likely overestimated outside of channels, given the spatial uniformity of high values, even on elevated road cells and watershed divides. For the higher depths considered (4 and 6 feet), synthetic flood probabilities have more distinct values between low-lying and high-lying areas, and align more closely to the latest 100-year and 500-year floodplain extents 27.
Synthetic flood probability overestimation can be explained because the three storm characteristics, i.e., duration, cumulative rainfall and peak intensity, are generated at the cell level independently, without considering the spatial autocorrelation existing in an actual storm. Future work will focus on additional characteristics that capture the spatial nature of precipitation events. These include the spatial autocorrelation of storm events of different duration, intensity and extent, the directions along which storms typically moves, and other spatio-temporal properties defined at the event-scale rather than cell-wise, e.g., rotation and velocity.
4.2 Computational efficiency
The computational complexity of the pipeline is overall significant. The cell-wise depth estimator costs higher computational resource than the universal models tested (Section 3.1), taking approximately 10 minutes and 100GB CPU memory to generate 100 precipitation-flood events. This performance is however competitive over HEC-RAS 2D, for which 15 found that the runtime for a single simulation was around 3 minutes, depending on the storm extent and duration. The obvious advantage of MaxFloodCast is the cell-model parallelization, which is not possible in HEC-RAS 2D. Future work should improve the efficiency of the computational process, especially in presence of additional relationships of spatial and hydrological nature that would enhance precipitation event generation.
4.3 Applications and future research
The Precipitation-Flood Depth Generative Pipeline was applied in Harris County due to the availability of an initial dataset of storms for training, as well as of a calibrated physics-based model that provided the data used to train a depth-estimator. The availability of historical storms and the resulting flood depth levels is necessary for the pipeline application in any study area, whether it is a prevalently natural, rural or urbanized region. Different model parametrization might be necessary to adapt to region-specific meteorological and hydrological conditions, especially when generating synthetic storms. Future research should determine what the appropriate scale is for the pipeline application, and if the scale difference between weather analysis (regional scale) and flood depth analysis (watershed scale) represents an issue for implementing the pipeline and producing accurate flood probability maps.
@@ 4.3 Applications and future research
Flood probability maps provide a spatial analysis of flood risk as a function of water depth. These maps emphasize the importance of depth-specific flood management strategies. This is crucial in regions where intricate flow dynamics and a flat ground surface significantly influence flood behavior. The high resolution of the maps also allows for practical applications, such as assessing the flood risk of buildings by integrating lowest floor elevation datasets, or identifying structures at risk of inundation a certain threshold. Additionally, these maps can evaluate transportation network vulnerability, as roads in areas with a high probability of exceeding 2 ft of inundation would become impassable, disrupting transportation and emergency response efforts. The non-binary nature of the maps further allows for comparative flood risk assessments across different inundation levels, enabling more nuanced and targeted flood mitigation strategies across diverse landscapes.
@@
Finally, this study also contributes to efforts to expand synthetic data for AI models developed for disaster resilience use cases. Both real and synthetic data support the development of robust AI models for various disaster resilience use cases ranging from risk prediction to situational awareness and impact assessment. Future studies can leverage GAN and other generative ML methods for augmenting data from occurred disasters and produce datasets suitable for training and testing novel ML models in multiple disaster resilience applications.
5 Conclusions
In this study, we proposed the Precipitation-Flood Depth Generative Pipeline, a novel machine learning framework for generating synthetic precipitation-flood events to construct high-resolution probabilistic flood maps. By integrating a Conditional Generative Adversarial Network (CTGAN) for synthetic precipitation generation with a specialized cell-wise depth estimator (MaxFloodCast V2), our approach addresses some of the limitations in traditional flood mapping methods, including data scarcity and computational costs associated with physics-based simulations. The results show the capability of the pipeline to produce flood probability maps that align with hydrological intuition and known flood-prone areas, highlighting its potential as a scalable tool for flood risk assessment.
Despite its advantages, certain limitations remain, such as potential inconsistencies in extreme precipitation cases and challenges in fully capturing spatial dependencies. Future research will focus on enhancing the integration of generative models with physics-based simulations, improving event-generation mechanisms to better capture localized precipitations, and further benchmarking against alternative deep learning models. Additionally, extending this framework to different geographical regions and incorporating real-time flood forecasting applications could expand its utility in disaster resilience and urban planning.
Acknowledgments
@@
The authors gratefully acknowledge the financial support provided by the National Science Foundation under the CRISP 2.0 Type 2 grant (grant number 1832662) and the Texas A&M X-Grant Presidential Excellence Fund. The perspectives, results, conclusions, and suggestions presented in this research are entirely those of the authors and do not necessarily represent the viewpoints of the funding organizations.
Code Availability